# Interpreting the Linear Structure of Vision-language Model Embedding Spaces

### v2

@@ Abstract
Vision-language models encode images and text in a joint space, minimizing the distance between corresponding image and text pairs. How are language and images organized in this joint space, and how do the models encode meaning and modality? To investigate this, we train and release sparse autoencoders (SAEs) on the embedding spaces of four vision-language models (CLIP, SigLIP, SigLIP2, and AIMv2). SAEs approximate model embeddings as sparse linear combinations of learned directions, or “concepts”. We find that, compared to other methods of linear feature learning, SAEs are better at reconstructing the real embeddings, while also able to retain the most sparsity. Retraining SAEs with different seeds or different data diet leads to two findings: the rare, specific concepts captured by the SAEs are liable to change drastically, but we also show that the key commonly-activating concepts extracted by SAEs are remarkably stable across runs. Interestingly, while most concepts are strongly unimodal in activation, we find they are not merely encoding modality per se. Many lie close to – but not entirely within – the subspace defining modality, suggesting that they encode cross-modal semantics despite their unimodal usage. To quantify this bridging behavior, we introduce the Bridge Score, a metric that identifies concept pairs which are both co-activated across aligned image-text inputs and geometrically aligned in the shared space. This reveals that even unimodal concepts can collaborate to support cross-modal integration. We release interactive demos of the SAEs for all models, allowing researchers to explore the organization of the concept spaces.
Overall, our findings uncover a sparse linear structure within VLM embedding spaces that is shaped by modality, yet stitched together through latent bridges—offering new insight into how multimodal meaning is constructed.
footnotetext: These authors contributed equally to this work.
@@ 1 Introduction
  • We develop VLM-Explore (https://vlm-concept-visualization.com/), an interactive visualization tool that lets researchers explore the linear concepts in a model. VLM-Explore helps visualize which concepts are important in achieving the model’s cross-modal spatial organization by visualizing our BridgeScore metric between pairs of concepts, quantifying how much they contribute to making a joint vision-language space (Section 4)
@@
  • We show that almost all concepts are single-modality concepts: they activate almost exclusively either for text or for image concepts, and this holds across all models that we try (Section 6)
  • However, we also show that this does not mean that the concepts lie within the separate cones of the image and text activations: across models, a large proportion of concepts are almost orthogonal (but not totally orthogonal) to the subspace that encodes modality, and therefore encode largely cross-modal dimensions of meaning (Section 6).
@@ 3 Methods: Sparse Autoencoders on Vision-Language Models
Figure 1: Sparsity–Expressivity trade-off. Pareto fronts for five dictionary learning methods applied to four vision-language models (CLIP, SigLIP, SigLIP 2, AIMv2). Each curve shows the trade-off between reconstruction quality ($R^{2}$ score) and sparsity level ($\ell_{0}$ norm of $\mathbf{Z}$). The three SAEs (TopK, JumpReLU, and BatchTopK) consistently achieve the best balance, with BatchTopK slightly dominating other sparse autoencoder variants.
Background and Notation.
We denote vectors by lowercase bold letters (e.g., $\bm{x}$) and matrices by uppercase bold letters (e.g., $\bm{X}$). We write $[n]$ to denote the index set $\{1,\dots,n\}$. The unit $\ell_{2}$-ball in $\mathbb{R}^{d}$ is defined as =$\mathcal{B}=\{\bm{x}\in\mathbb{R}^{d}\mid\|\bm{x}\|_{2}\leq 1\}$. We consider a general multimodal representation learning setting in which a vision-language model maps image or text inputs $\bm{x}\in\mathcal{X}$ into a shared representation space $\mathcal{A}\subseteq\mathbb{R}^{d}$. A collection of $n$ such embeddings is stored in a matrix $\bm{A}\in\mathbb{R}^{n\times d}$. Dictionary Learning seeks to approximate $\bm{A}$ as a linear combination of concept vectors from a dictionary $\bm{D}\in\mathbb{R}^{c\times d}$, using a sparse code matrix $\bm{Z}\in\mathbb{R}^{n\times c}$:
@@
Expressivity vs Sparsity.
An ideal dictionary learning method should produce representations that are both expressive and sparse. In practice, this means that model embeddings should activate only a small number of meaningful concepts, indicating that the learned dictionary is both specific and semantically aligned with the data. We apply five dictionary learning methods to four models (CLIP, SigLIP and SigLIP2, which are encoder models, and AIMv2 which is an autregressive vision-language model). The activation matrix $\bm{A}$ consists of 600,000 normalized embeddings from passing the COCO dataset (Lin et al., 2014) through the models, with each $\bm{A}_{i}\in\mathcal{B}$, the unit $\ell_{2}$-ball.
We evaluate all SAEs on their trade-off between sparsity – measured as the $\ell_{0}$ norm of $\bm{Z}$, i.e., the number of non-zero entries per code vector – and reconstruction quality, quantified by the $R^{2}$ score between the original activations $\bm{A}$ and their reconstructions $\bm{Z}\bm{D}$. This trade-off captures how well the learned codes retain information while encouraging interpretability through sparsity. Results, shown in Figure 1, indicate that all four SAEs variants outperform Semi-NMF at fixed sparsity levels, with BatchTopK achieving slightly better performance overall. In practice, we also find that training TopK and BatchTopK SAEs is more stable and requires less hyperparameter tuning. For these reasons, we use the BatchTopK SAE as our default model in all subsequent experiments.
@@ 3.1 Metrics: The Geometry and Statistics of Learned Concepts
To assess the interpretability and quality of the learned representations, we introduce four complementary metrics that go beyond the standard sparsity–reconstruction trade-off: stability, energy, modality score, and bridge score. Each of these measures captures a different axis of representational information to better understand the representation of VLM using the SAEs.
@@
Stability quantifies the consistency of learned dictionaries under different training runs with random initialization. If two independently trained SAEs on the same model produce significantly different dictionaries, it suggests that the learned concepts are not robust, which weakens their interpretability. Following Fel et al. (2025) and Spielman et al. (2012), we compute the stability between two dictionaries $\bm{D},\bm{D}^{\prime}\in\mathbb{R}^{c\times d}$ by optimally aligning their rows using the Hungarian algorithm, which solves for a permutation matrix $\bm{P}\in\mathcal{P}(c)$ that maximizes the total similarity between matched concept vectors:
@@
Energy.
While stability measures whether concepts are consistently learned, it does not tell us which concepts matter most in practice. To better understand which concepts are actually used by the model, we turn to the energy of each concept. Energy captures the average activation strength of a concept across a dataset, revealing which parts of the dictionary are consistently involved in reconstructions. Formally, for concept $i$, we define its energy as: $\text{Energy}_{i}=\mathbb{E}_{\bm{z}}(z_{i})$. Energy provides a statistical view over the dataset, revealing where attention should be focused for interpreting the learned representation.
@@
A modality score near $1$ indicates that the concept fires only image inputs, while a score near 00 suggests text dominance. A value around $\frac{1}{2}$ indicates a multi-modal concept.
@@
The resulting matrix $\mathbf{B}$ reveals how concepts jointly operate across modalities. A large value at position $\mathbf{B}_{i,j}$ indicates that concepts $i$ and $j$ frequently co-activate and encode similar directions in the shared embedding space. These concept pairs act as semantic bridges, facilitating the alignment of image and text embeddings.
In this sense, the bridge matrix provides insight into why the model succeeds at cross-modal alignment. It highlights not just which concepts are shared, but how they collaborate structurally and statistically to support the model’s semantic integration of vision and language.
4 VLM-Explore: An interactive, functional concept explorer
Our first contribution is the development of VLM-Explore, an interactive visualization tool designed to facilitate the exploration and analysis of concept representations extracted from VLM using SAEs. This platform provides a functional interface for examining the internal structure and organization of learned concepts within the joint vision-language embedding space. The core view of VLM-Explore presents a two-dimensional UMAP projection of the SAE dictionary $\bm{D}$, where each point corresponds to a learned concept vector. The visual encoding incorporates both statistical and semantic information: the size of each point represents the energy of the concept, and the color of each point encodes its modality score (Equation 5). This visualization enables intuitive, high-level assessment of how concepts are distributed with respect to modality and influence within the learned representation.
Upon selecting a concept, the interface displays its top activating samples from the COCO dataset, thereby offering concrete examples of the semantic attributes captured by the concept. A central feature of VLM-Explore is its “co-occurrence” view, which renders edges from the selected concept to the concepts with the highest bridge score. The thickness of each edge reflects the magnitude of this score, thereby emphasizing concept pairs that contribute substantially to cross-modal alignment. This interactive graph-based representation offers insight into the latent relational structure of the concept space, revealing both local interactions and global patterns of semantic organization.
By supporting the identification of dominant, modality-specific, and cross-modal bridging concepts, VLM-Explore serves as a diagnostic and interpretive tool for understanding how VLMs organize semantic content across modalities. Concept pairs with high bridge scores, in particular, can be viewed as structural components that facilitate the alignment objective implicitly optimized by the underlying VLM. We release VLM-Explore in conjunction with all trained SAE dictionaries for the four models analyzed in this work—CLIP, SigLIP, SigLIP2, and AIMv2—offering an open platform for further research into the compositional and linear structure of multimodal embedding spaces.
Figure 2: Stability vs. Energy. (Left) Energy is concentrated on a few concepts, with 512 concepts getting 99% of the energy, and the other 3,500 concepts only sharing 1% of the total coefficient weight. (Center) When training with 5 different random seeds, the stability for all 4,096 concepts is low – 0.16 (Right) When we weight concepts by energy, we can see that the concepts that are used often in reconstruction are actually stable, with stability of 0.92 if we take the top 512 energetic concepts from each run. The instability comes from concepts that are rarely if ever actually used.
@@ 5 Analysis 1: Are We Finding Consistent Model Features?
Before analyzing the semantic content of the concepts extracted by our SAEs (Section 6), we first examine whether these concepts are consistent—both across training runs with different random seeds, and across variations in the SAE training dataset. We study stability across: (i) random seeds—to test sensitivity to training stochasticity, and (ii) data modality mixtures—to test dependence on the distribution of input activations.
Results: Stability Across Different Seeds
@@
We train five SAEs with different random seeds on the same SigLIP2 activations, each using $4096$ concepts, and compute the average pairwise stability (Equation 4) between them. As shown in Figure 2 (center), at first glance, overall stability is low: the mean similarity between dictionaries is just $0.16$, suggesting high sensitivity to random initialization. However, this apparent instability is misleading. On the left, see that reconstruction energy is heavily concentrated in a small subset of concepts: the top $512$ concepts account for approximately $99\%$ of the total activation mass. To isolate the functionally relevant features, we recompute stability using only the top-$k$ most energetic concepts from each run. We find that high-energy concepts are remarkably stable, with near-perfect alignment across seeds (Figure 2, right). This reveals a key paradox:
Finding: Without conditioning on energy, SAEs concepts appear unstable; but when viewed in conjunction with their associated energy, a consistent and robust set of model features emerges.
Figure 3: UMAP visualization of the SAE concept spaces under different image–text data mixtures. Color indicates the modality score of each concept. The top row displays all concepts with equal size; the bottom row scales point size by concept energy. While the dominant modality in the training data strongly influences the number of concepts recovered (top row), the most energetic concepts (bottom row) remain relatively stable across mixtures. A similar set of high-energy text and image concepts is recovered regardless of the input distribution. Additional results for other models are provided in Appendix A, Figure 9.
The instability arises almost entirely from the low-energy tail of the dictionary—concepts that contribute little to the reconstruction objective. We conclude that SAEs trained with different seeds reliably recover a core set of functional, high-usage concepts.
Results: Stability Across Different Training Data Mixtures
@@
The analysis above addresses random variation, but it remains unclear whether the recovered concepts reflect structure in the model or simply the statistics of the dataset used to train the SAE. To test this, we vary the modality mixture of the SAE training activations: how many come from image inputs vs. text inputs. We train SAEs on seven data compositions, ranging from heavily image-dominant (5:1 image:text) to heavily text-dominant (1:5). Results are shown in Figure 3. As expected, changing the data mixture shifts the overall modality distribution of the learned concepts. However, when weighting by energy (bottom row), the most energetic concepts remain stable across all mixtures. High-energy directions consistently appear regardless of the training distribution, and exhibit similar activation patterns.
These analyses show that while SAE dictionaries may appear unstable, this instability is concentrated in low-energy concepts. The high-energy subset—responsible for nearly all reconstruction—is both robust to random initialization and is also not simply a reflection of the data. Crucially, this consistency only holds when conditioning on energy, highlighting that stability emerges from how the model uses the learned concepts.
@@
Our evaluations point to two main findings. First, the important (i.e., high-energy) concepts in our SAEs are stable: they remain consistent across different seeds and are largely unaffected by changes in the image-text data mixture. Second, the low-energy concepts are highly unstable, with different SAEs discovering widely different, rarely-used features. Does this mean that low-energy concepts are useless? Not necessarily. Leveraging VLM-Explore (Section 4) we observe that many low-energy concepts correspond to coherent but rare patterns, such as “yaks” or “valentines day hearts.” These concepts may be semantically meaningful, but their contribution to reconstruction is small, and thus SAEs do not consistently recover them. By contrast, high-energy concepts typically capture more concepts like “red” or “two women” and are consistently recovered because they support a wide range of reconstructions.
These findings guide the rest of this paper. They suggest that SAEs can be used as reliable tools for interpreting VLM embeddings—but only when conditioned on energy. Concepts with high energy form a stable and informative basis; concepts with low energy may still be meaningful, but should be treated with caution due to their instability.

### v3

@@ Abstract
Vision-language models encode images and text in a joint space, minimizing the distance between corresponding image and text pairs. How are language and images organized in this joint space, and how do the models encode meaning and modality? To investigate this, we train and release sparse autoencoders (SAEs) on the embedding spaces of four vision-language models (CLIP, SigLIP, SigLIP2, and AIMv2). SAEs approximate model embeddings as sparse linear combinations of learned directions, or “concepts”. We find that, compared to other methods of linear feature learning, SAEs are better at reconstructing the real embeddings, while also able to retain the most sparsity. Retraining SAEs with different seeds or different data diet leads to two findings: the rare, specific concepts captured by the SAEs are liable to change drastically, but we also show that commonly-activating concepts are remarkably stable across runs. Interestingly, while most concepts activate primarily for one modality, we find they are not merely encoding modality per se. Many are almost orthogonal to the subspace that defines modality, and the concept directions do not function as good modality classifiers, suggesting that they encode cross-modal semantics. To quantify this bridging behavior, we introduce the Bridge Score, a metric that identifies concept pairs which are both co-activated across aligned image-text inputs and geometrically aligned in the shared space. This reveals that even single-modality concepts can collaborate to support cross-modal integration. We release interactive demos of the SAEs for all models, allowing researchers to explore the organization of the concept spaces. Overall, our findings uncover a sparse linear structure within VLM embedding spaces that is shaped by modality, yet stitched together through latent bridges—offering new insight into how multimodal meaning is constructed.
footnotetext: * These authors contributed equally to this work.
@@ 1 Introduction
  • We show that almost all concepts are single-modality concepts: they activate almost exclusively either for text or for image concepts, and this holds across all models that we try (Section 5.1)
  • However, we also show that this does not mean that the concepts lie within the separate cones of the image and text activations: across models, a large proportion of concepts are almost orthogonal (but not totally orthogonal) to the subspace that encodes modality, and therefore encode largely cross-modal dimensions of meaning (Section 5.2).
  • We develop VLM-Explore (https://vlm-concept-visualization.com/), an interactive visualization tool that lets researchers explore the linear concepts in a model, and how they connect modalities (Section 6).
@@ 3 Methods: Sparse Autoencoders on Vision-Language Models
Figure 1: Selecting a sparse dictionary learning method: the Expressivity-Sparsity trade-off. Pareto fronts for five dictionary learning methods applied to four vision-language models (CLIP, SigLIP, SigLIP 2, AIMv2). Each curve shows the trade-off between reconstruction quality ($R^{2}$ score) and sparsity level ($\ell_{0}$ norm of $\mathbf{Z}$). The three SAEs (TopK, JumpReLU, and BatchTopK) consistently achieve the best balance, with BatchTopK slightly dominating other sparse autoencoder variants.
3.1 Background and Notation.
We denote vectors by lowercase bold letters (e.g., $\bm{x}$) and matrices by uppercase bold letters (e.g., $\bm{X}$). We write $[n]$ to denote the index set $\{1,\dots,n\}$. The unit $\ell_{2}$-ball in $\mathbb{R}^{d}$ is defined as $\mathcal{B}=\{\bm{x}\in\mathbb{R}^{d}\mid\|\bm{x}\|_{2}\leq 1\}$. We consider a general multimodal representation learning setting in which a vision-language model maps image or text inputs $\bm{x}\in\mathcal{X}$ into a shared representation space $\mathcal{A}\subseteq\mathbb{R}^{d}$. A collection of $n$ such embeddings is stored in a matrix $\bm{A}\in\mathbb{R}^{n\times d}$. Dictionary Learning seeks to approximate $\bm{A}$ as a linear combination of concept vectors from a dictionary $\bm{D}\in\mathbb{R}^{c\times d}$, using a sparse code matrix $\bm{Z}\in\mathbb{R}^{n\times c}$:
@@
3.2 Selecting a sparse dictionary method: Expressivity vs sparsity
An ideal dictionary learning method should produce representations that are both expressive and sparse. In practice, this means that model embeddings should activate only a small number of meaningful concepts, indicating that the learned dictionary is both specific and semantically aligned with the data. In our first set of preliminary experiments, we compare five methods on this trade-off, and present our results in Figure 1.
Dictionary learning methods:
We test five different dictionary learning methods: Semi-Nonnegative Matrix Factorization (Semi-NMF) (Parekh et al., 2025; Lee & Seung, 1999), Sparse Autoencoders (SAEs), Top-K SAEs (Gao et al., 2025), JumpReLU SAEs (Rajamanoharan et al., 2024), and BatchTopK SAEs (Bussmann et al., 2024), and train using Overcompletehttps://github.com/KempnerInstitute/overcomplete.
Models:
We run experiments on four models: CLIP (Radford et al., 2021), SigLIP (Zhai et al., 2023), and SigLIP2 (Tschannen et al., 2025), which are encoder models, and AIMv2 (El-Nouby et al., 2024; Fini et al., 2024), which is an autoregressive vision-language model
Data:
We train all dictionary learning methods on reconstructing the activation matrix $\bm{A}$, consisting of 600,000 normalized embeddings from passing the COCO dataset (Lin et al., 2014) through the models, with each $\bm{A}_{i}\in\mathcal{B}$, the unit $\ell_{2}$-ball
Results:
We evaluate the expressivity and sparsity of all methods, and present our results in Figure 1. We measure sparsity as the number of non-zero entries in the code vector (i.e., the $\ell_{0}$ of $Z$). For different levels of sparsity, we plot the reconstruction quality (expressivity): the $R^{2}$ score between the original activations $\bm{A}$ and their reconstructions $\bm{Z}\bm{D}$. The trade-off captures how well the learned codes retain information while encouraging interpretability through sparsity.
Our results indicate that all four SAE variants outperform Semi-NMF at fixed sparsity levels, with BatchTopK achieving slightly better performance overall. In practice, we also find that training TopK and BatchTopK SAEs is more stable and requires less hyperparameter tuning. For these reasons, we use the BatchTopK SAEs for our subsequent analysis experiments.
@@ 3.3 Metrics: The geometry and statistics of learned concepts
To assess the interpretability and quality of the learned representations, we introduce four complementary metrics that go beyond the standard sparsity–reconstruction trade-off: energy, stability, modality score, and bridge score. Each of these measures captures a different axis of representational information:
Energy.
Energy is a measure of how often each concept is used: its average activation strength across a dataset. Formally, for a concept $i$, we define its energy as: $\text{Energy}_{i}=\mathbb{E}_{\bm{z}}(z_{i})$. Energy provides a statistical view over the dataset, revealing where attention should be focused for interpreting the learned representation.
@@
Stability quantifies how consistent the learned dictionaries are between two runs. If two independently trained SAEs on the same model produce significantly different dictionaries, it suggests that the learned concepts are not robust, which weakens their interpretability. Following Fel et al. (2025) and Spielman et al. (2012), we compute the stability between two dictionaries $\bm{D},\bm{D}^{\prime}\in\mathbb{R}^{c\times d}$ by optimally aligning their rows using the Hungarian algorithm, which solves for a permutation matrix $\bm{P}\in\mathcal{P}(c)$ that maximizes the total similarity between matched concept vectors:
@@
A modality score near $1$ indicates that the concept fires only image inputs, while a score near 0 suggests text dominance. A value around $\frac{1}{2}$ indicates a multi-modal concept.
Figure 2: The concepts that use most of the energy are stable Here we compare the relationship between energy and stability: (Left) Energy is concentrated on a few concepts, with 512 concepts getting 99% of the energy, and the other 3,500 concepts only appearing in 1% of the total coefficient weight. (Center) When training with 5 different random seeds, the stability when we consider all 4,096 concepts is low – 0.16 (Right) When we weight concepts by energy, we can see that the concepts that are used often in reconstruction are actually stable, with stability of 0.92 if we take the top 512 energetic concepts from each run. The instability comes from concepts that are rarely if ever actually used.
@@
The resulting matrix $\mathbf{B}$ reveals how concepts jointly operate across modalities. In this sense, the bridge matrix provides insight into why the model succeeds at cross-modal alignment. It highlights not just which concepts are shared, but how they collaborate structurally and statistically to support the model’s semantic integration of vision and language.
Figure 3: The geometry of high-energy concepts is stable across data mixtures. Top: each concept has equal size, Bottom: size is dependent on energy. UMAP visualization of the SAE concept spaces under different image–text data mixtures. Color indicates the modality score of each concept. While the dominant modality in the training data strongly influences how many concepts are recovered per modality (top row), the most energetic concepts (bottom row) remain relatively stable across mixtures. The set of high-energy text and image concepts remains relatively consistent regardless of the input distribution. Additional results for the other three models are provided in Appendix A, Figure 9.
@@ 4 Analysis 1: Are We Finding Consistent Model Features?
Before analyzing the geometry and semantic content of the concepts extracted by our SAEs (Section 5), we first examine whether these concepts are consistent and stable: do we get similar concepts across different conditions? We study stability across: (i) random seeds—to test sensitivity to training stochasticity, and (ii) data modality mixtures—to test dependence on the distribution of input activations.
4.1 The concepts that use most of the energy are stable
@@
We train five SAEs with different random seeds on the same SigLIP2 activations, each using $4096$ concepts, and compute the average pairwise stability (Equation 4) between them. As shown in Figure 2 (center), at first glance, overall stability is low: the mean similarity between dictionaries is just $0.16$, suggesting high sensitivity to random initialization. However, this apparent instability is misleading. On the left, see that reconstruction energy is heavily concentrated in a small subset of concepts: the top $512$ concepts account for approximately $99\%$ of the total activation mass. To isolate the functionally relevant features, we recompute stability using only the top-$k$ most energetic concepts from each run. We find that high-energy concepts are remarkably stable, with near-perfect alignment across seeds (Figure 2, right). The instability arises almost entirely from the low-energy tail of the dictionary—concepts that contribute little to the reconstruction objective. We conclude that SAEs trained with different seeds reliably recover a core set of functional, high-usage concepts.
4.2 The geometry of high-energy concepts is stable across data mixtures
@@
Though we have seen that the high-use concepts retrieved by the SAE are stable across random initializations, there is still the question of how much the recovered concepts are influenced by the data used to train the SAEs. Do our SAE dictionaries reflect the structure in the model or simply the statistics of the dataset used to train the SAE? To test this, we train SAEs varying the modality mixture of the training data: how many of the activations we use to train the SAE come from image inputs vs. text inputs. We train SAEs on seven data compositions, ranging from heavily image-dominant (5:1 image:text) to heavily text-dominant (1:5). Results are shown in Figure 3, where we contrast the UMAP geometry of the learned concepts concepts (top) with the geometry of the concepts when the size of each point is weighted by the concept’s energy. As expected, changing the data mixture shifts the overall modality distribution of the learned concepts: the text-heavy SAEs learn more text concepts (blue points) and the image-heavy SAEs learn more image concepts. However, when weighting by energy (bottom row), the geometry of the most energetic concepts remain stable across all mixtures: the visualizations look remarkably consistent across data mixtures. High-energy directions consistently appear regardless of the training distribution, and exhibit similar activation patterns.
Figure 4: Most concepts are single-modality Histograms of the modality scores (top) and the modality scores weighted by energy (bottom) of every concept in the SAEs. On every model, the modes are the two extremes: concepts either activated only by text or only by image. Weighting by energy makes this much more prominent, showcasing that almost all of the reconstruction coefficients go to concepts that are single-modality.
@@
Our evaluations show that different SAEs discover widely different low-energy features. Does Does this mean that low-energy concepts are useless? Not necessarily. Leveraging VLM-Explore (Section 6) we observe that many low-energy concepts correspond to coherent but rare patterns, such as “yaks” or “valentines day hearts.” These concepts may be semantically meaningful, but their contribution to reconstruction is small, and thus SAEs do not consistently recover them. By contrast, high-energy concepts typically capture more concepts like “red” or “two women” and are consistently recovered because they support a wide range of reconstructions. We provide some illustrative examples of concepts in Appendix B.
These analyses show that while SAE dictionaries may appear unstable, this instability is concentrated in low-energy concepts. The high-energy subset—responsible for nearly all reconstruction—is both robust to random initialization and is also not simply a reflection of the data. These findings guide the rest of this paper. They suggest that SAEs can be used as reliable tools for interpreting VLM embeddings—but only when conditioned on energy. Concepts with high energy form a stable and informative basis; concepts with low energy may still be meaningful, but should be treated with caution due to their instability.